Suppose Brittany has the following demand curve for oranges:

Q = 90 − 3P

  1. If oranges cost $2, what is her own-price elasticity of demand for oranges?
Answer #1

Q= 90-3P

(Change in Q/ Change in P)= -3

P= $2

Q= 90-3(2)= 90-6 = 84 units

Own -price elasticity of demand for oranges =(Change in Q/ Change in P) (P/Q)

= (-3)(2/84) = -0.07